# On the inadequacy of nudging data assimilation algorithms for non-dissipative systems

### v2

@@ 3.2.1. Classical KdV equation
We note that the simulations presented here indicate that nudging is ill-suited to extremely capture high oscillatory solutions to Equation˜19, but it does not necessarily prohibit other methods of data assimilation from recovering the reference solution. Unlike in Figure˜5, where the observed data is identically zero, it is not impossible in theory for data assimilation to tell the difference between two solutions initialized of the form Equation˜29. That is, unlike the case of Figure˜5 if one initializes two solutions $u_{1}$ and $u_{2}$ of Equation˜19 with initial data prescribed by Equation˜29 with separate $k_{0}$ values for $u_{1}$ and $u_{2}$, we have that $\mathinner{\!\left\lVert P_{M}(u_{1}-u_{2})\right\rVert}_{L^{2}}\neq 0$ for $t>0$, and to the best of our knowledge one cannot readily construct two such solutions of this form that always agree on every observed wavemode for all time. This is due to the fact that dispersive term essentially causes the coefficients of the wavemodes to rotate in the complex plane around the origin with rotation speed $\delta^{2}k^{3}$. As each wavemode is rotating at a different speed one cannot rescale and shift momentum across wavemodes in a straightforward way such that the momentum backscattering due to the nonlinearity on the observed modes still agree at each timestep to produce a counterexample similar to the one present in Figures˜5 and 4.
We note that all of these cases where nudging fails are specific to the undamped and unforced, classical KdV Equation˜19, which is not a dissipative dynamical system. The addition of damping will result in the energy being drained from these high oscillations, while the forcing injects energy into specified low wavemodes, resulting in nontrivial long time dynamics that eliminate these sorts of extremely high frequency oscillations, see Section˜3.2.2 below.
@@ 3.2.2. Damped and Driven KdV
In [53] it was shown analytically, that given enough observed Fourier modes and large enough $\mu$, that the nudged solution converges to the reference solution in $L^{2}$ exponentially fast in time with exponent $\frac{\gamma}{4}$ for arbitrary choice of initial data $v^{in}\in H^{2}_{\text{per}}$ with spatial mean zero for Equation˜33.
@@
We now verify computationally that nudging successfully drives the assimilated solution to the reference solution exponentially fast in time. Depicted in Figure˜9 we can see nudging effectively recovers a reference solutions initialized as in Equation˜29 exponentially fast in time. Here we utilize $\mu=100$ and $\gamma=0.1$ and we note that the exponential rate of convergence is given by small perturbations from $\gamma$. Here we initialize the reference solution as in Equation˜29 with $k_{0}=12$, that is, we selected the initial profile $u^{in}=\cos(12\cdot\pi\cdot x)$. For our driving force, we utilized $f=\cos(8\cdot\pi\cdot x)$. We note that and we observed only the first 10 modes of the solution. It is unsurprising that convergence happens at the exponential rate proportional to the damping coefficient, $\gamma$, as this coincides with the results of [53]. We note that the exponential convergence is not entirely independent of $\mu$, the oscillatory component at initial times and on the observed modes is particularly sensitive to $\mu$ values, although this effect quickly fades as the error begins to decay at rate $\gamma$. Moreover, as one sees from the spectrum plots in Figure˜9 the primary barrier for convergence are the spikes in undetected frequencies, therefore it is unsurprising that convergence to the reference solution happens at exactly the rate, $\gamma$, at which momentum is drained from these modes. Indeed, this behavior is typical of simulations of nudging methods, as the unobserved modes rely on the dissipation in the system in order to synchronize. The overall decay in error can typically be decomposed into two distinct phases, an initial rapid decay in error that is highly dependent on $\mu$, and a slower, but still exponentially fast in time, rate of convergence primarily driven by the dissipation in the system. For an in-depth study of the role of the particular value of $\mu$ see, e.g., [15].
@@
We note that the AOT algorithm successfully recovers the reference solution, $u$, to Equation˜32 even when $u$ is initialized as in the counterexamples from the previous section Equations˜29 and 31 given enough observed modes. We also point out, that in Figure˜9 we see some behavior in the observed modes due to the presence of the body force. We could disable the body force by setting it to 0, however in that case data assimilation would not be required at all. This is because the damping term will drive any initial condition to 0, exponentially fast at the rate $\gamma$ with or without any form of data assimilation being applied.
@@ 4.1. Analytical Results
We now move to extend our results to the incompressible Euler equations on a 2D periodic domain. The equations are given as follows:
$\begin{split}\boldsymbol{u}_{t}+\boldsymbol{u}\cdot\nabla\boldsymbol{u}&=-\nabla p,\\ \divergence\boldsymbol{u}&=0.\end{split}$
@@
$\begin{split}\boldsymbol{v}_{t}+\boldsymbol{v}\cdot\nabla\boldsymbol{v}&=-\nabla q+\mu P_{M}(\boldsymbol{u}-\boldsymbol{v}),\\ \divergence\boldsymbol{v}&=0.\end{split}$
Here $\boldsymbol{v}(x,t)=(v_{1}(\boldsymbol{x},t),v_{2}(\boldsymbol{x},t))$ is the velocity field, $q$ is the pressure, $\mu$ is the nudging parameter, and $M$ is the number of observed modes in the reference solution $u$ of Section˜4.1. As we are observing a 2D solution, $u$ of the system Section˜4.1, $P_{M}$ denotes the observations of all Fourier modes $\boldsymbol{k}\in\mathbb{R}^{2}$ such that $\mathinner{\!\left\lvert\boldsymbol{k}\right\rvert}\leq M$.
@@
$(\boldsymbol{u},\boldsymbol{v})=\sum_{i=1}^{2}\int_{\Omega}u_{i}v_{i}d{\boldsymbol{x}}\quad\text{ and }\quad((\boldsymbol{u},\boldsymbol{v}))=\sum_{i,j=1}^{2}\int_{\Omega}\partialderivative{u_{i}}{x_{j}}\partialderivative{v_{i}}{x_{j}}d{\boldsymbol{x}},$ (35)
which are associated with the norms $\mathinner{\!\left\lVert\boldsymbol{u}\right\rVert}_{L^{2}}=\mathinner{\!\left\lVert\boldsymbol{u}\right\rVert}_{L^{2}}=(\boldsymbol{u},\boldsymbol{u})^{1/2}$ and $\mathinner{\!\left\lVert\boldsymbol{v}\right\rVert}_{V}=((\boldsymbol{u},\boldsymbol{u}))^{1/2}$. We note that the $V$ norm is the $H^{1}$ seminorm, which we use as the full norm for $V$ which is justified by the Poincaré inequality. We denote by $V^{s}$ the equivalent Sobolev space, given by $V^{s}=\overline{\mathcal{V}}^{V^{s}}$, where the $V^{s}$ norm is given by the $H^{s}$ seminorm.
We now apply the Leray-Helmholtz projector, $P_{\sigma}$, to Sections˜4.1 and 4.1 to formally obtain the following equations for $\boldsymbol{u}$ and $\boldsymbol{v}$, respectively
@@
Here we formally denote the nonlinear operator, $B(\boldsymbol{a},\boldsymbol{b})\mathrel{\mathop{\ordinarycolon}}=P_{\sigma}(\boldsymbol{a}\cdot\gradient\boldsymbol{b})$.
We recall the following theorem regarding global well-posedness of solutions to Section˜4.1
Theorem 4.1 (Global well-posedness of Section˜4.1 (see e.g. [64])).
Given $\boldsymbol{u}^{in}\in V^{s}(\mathbb{T}^{2})$ with $s\geq 3$, then Section˜4.1 is globally well-posed with solution
@@
We now move to restate our propositions from Section˜3.2.1 in this context.
@@
Let $k\in\mathbb{Z}\setminus\{0\}$ and $\phi\in L^{2}_{\text{per}}$ with $\phi$ having spatial mean zero. The following hold true:
  1. (i)
    $\phi(\boldsymbol{x}+\frac{L}{k}\boldsymbol{e_{i}})=\phi(\boldsymbol{x})$ holds for a.e. $x\in\mathbb{R}^{2}$ for $i=1,2$ if and only if
    $\phi(\boldsymbol{x})=\sum_{\boldsymbol{n}\in k\mathbb{Z}^{2}}\hat{\phi}_{\boldsymbol{n}}e^{i\frac{2\pi}{L}\boldsymbol{n}\cdot\boldsymbol{x}}.$ (38)
  2. (ii)
    Let $M$ be a positive integer and let $k$ be an integer satisfying $\mathinner{\!\left\lvert k\right\rvert}\geq M$. Suppose $\phi(\boldsymbol{x}+\frac{L}{k}\boldsymbol{e}_{i})=\phi(\boldsymbol{x})$ for a.e. $x\in\mathbb{R}^{2}$ for $i=1,2$, then
    $P_{M}\phi=0$ (39)
@@
Suppose $k\in\mathbb{Z}\setminus\{0\}$ and $\boldsymbol{u}^{in}\in V^{3}$ such that $\boldsymbol{u}^{in}(\boldsymbol{x}+\frac{L}{k}\boldsymbol{e}_{i})=\boldsymbol{u}^{in}(x)$ for all $x\in\mathbb{R}$ and each $i=1,2$. Additionally suppose that $\boldsymbol{u}^{in}$ has spatial mean zero. Let $u(\boldsymbol{x},t)$ be the corresponding solution of Section˜4.1, subject to periodic boundary conditions with period $L$ in both the horizontal and vertical directions.
Then $u(\boldsymbol{x}+\frac{L}{k}\boldsymbol{e}_{i},t)=u(x,t)$ for a.e. $x\in\mathbb{R}$ and all $t\in\mathbb{R}$.
Proof.
Let $\boldsymbol{v}(\boldsymbol{x},t)=u(\boldsymbol{x}+\frac{L}{k}\boldsymbol{e}_{i},t)$ for a fixed $i=1$ or $2$ and notice that
$\boldsymbol{v}(\boldsymbol{x},0)=\boldsymbol{v}^{in}(\boldsymbol{x})=\boldsymbol{u}^{in}(\boldsymbol{x}+\frac{L}{k}\boldsymbol{e}_{i},0)=\boldsymbol{u}^{in}(\boldsymbol{x}).$
Thus we see that $\boldsymbol{v}(\boldsymbol{x},t)$ is also a solution of Section˜4.1. As $\boldsymbol{v}(\boldsymbol{x},0)=\boldsymbol{u}^{in}(\boldsymbol{x})$, then by the uniqueness of solutions to Section˜4.1 we obtain that $\boldsymbol{v}(\boldsymbol{x},t)=\boldsymbol{u}(\boldsymbol{x},t)$, hence we have that $\boldsymbol{u}(\boldsymbol{x}+\frac{L}{k}\boldsymbol{e}_{i},t)=\boldsymbol{u}(\boldsymbol{x},t)$. ∎
@@
Under the conditions of Proposition˜3.3 the space of functions periodic with respect to the lattice $k\mathbb{Z}^{2}$ is invariant under the solution of Section˜4.1.
Proposition 4.5.
Suppose $k\in\mathbb{Z}\setminus\{0\}$ is given and suppose $\boldsymbol{u}^{in}_{j}\in V^{3}$, for $j=1,2$, satisfying
$\boldsymbol{u}^{in}_{j}(\boldsymbol{x}+\frac{L}{k}\boldsymbol{e}_{i})=\boldsymbol{u}_{j}^{in}(\boldsymbol{x})$ (40)
for a.e. $\boldsymbol{x}\in\mathbb{R}^{2}$, $i=1,2$, and $\boldsymbol{u}_{j}^{in}$ has spatial mean zero. Moreover, assume $\mathinner{\!\left\lVert\boldsymbol{u}^{in}_{1}\right\rVert}_{L^{2}}\neq\mathinner{\!\left\lVert\boldsymbol{u}^{in}_{2}\right\rVert}_{L^{2}}$ and let $\boldsymbol{u}_{j}(x,t)$ be the corresponding solutions of Section˜4.1. Then
  1. (i)
    $\boldsymbol{u}_{j}(\boldsymbol{x}+\frac{L}{k}\boldsymbol{e}_{i},t)=\boldsymbol{u}_{j}(\boldsymbol{x},t)$ for a.e. $\boldsymbol{x}\in\mathbb{R}$, $i=1,2$, and for $j=1,2$ and all $t\in\mathbb{R}^{2}$.
  2. (ii)
    $P_{M}\boldsymbol{u}_{1}(\cdot,t)=P_{M}\boldsymbol{u}_{2}(\cdot,t)$ for any positive integer $M<\mathinner{\!\left\lvert k\right\rvert}$, and every $t\in\mathbb{R}^{2}$.
  3. (iii)
    $\limsup\limits_{t\to\infty}\mathinner{\!\left\lVert\boldsymbol{u}_{1}(\cdot,t)-\boldsymbol{u}_{2}(\cdot,t)\right\rVert}_{L^{2}}>0.$
Now, using these propositions, we show that Section˜4.1 does not have the finitely many determining modes property.
Theorem 4.6.
The incompressible Euler equations Section˜4.1 subject to periodic boundary conditions on $\mathbb{T}^{2}$ does not possess the finitely many determining modes property in $H$.
Corollary 4.7.
For a fixed number of observed Fourier modes, $M$, there exists solutions $\boldsymbol{u}(\boldsymbol{x},t)$ of Section˜4.1 that cannot be recovered by Section˜4.1 with arbitrary initial condition $\boldsymbol{v}^{in}$.
Proof.
Let $M\in\mathbb{N}$ be given and let $\boldsymbol{u}(\boldsymbol{x},t)$ be any nontrivial solution to Section˜4.1 that is periodic with respect to the lattice $k\mathbb{Z}^{2}$ for $k\in\mathbb{Z}$ given such that $\mathinner{\!\left\lvert k\right\rvert}>M$.
Now, as $\boldsymbol{v}^{in}$ is arbitrary we can specify $\boldsymbol{v}^{in}\neq\boldsymbol{u}^{in}$ with $\boldsymbol{v}^{in}$ also being periodic with respect to the same lattice as $\boldsymbol{u}$. We will additionally assume that $\mathinner{\!\left\lVert\boldsymbol{u}^{in}\right\rVert}_{H}\neq\mathinner{\!\left\lVert\boldsymbol{v}^{in}\right\rVert}_{H}$.
We note that, by construction, the nudging term $\mu P_{\sigma}P_{M}(\boldsymbol{u}-\boldsymbol{v})$ drops out entirely in Equation˜37, and thus one can easily verify that $\boldsymbol{v}(\boldsymbol{x},t)$ is in fact a solution to Section˜4.1 corresponding to the initial data $\boldsymbol{v}^{in}$. The result immediately follows from the fact that the $H$ norm is a conserved quantity for solutions to Section˜4.1. ∎

### v3

@@ 3.2.1. Classical KdV equation
The simulations presented here indicate that nudging is ill-suited to capture extremely high oscillatory solutions to Equation˜19, but it does not necessarily prohibit other methods of data assimilation from recovering the reference solution. Unlike in Figure˜5, where the observed data is identically zero, it is not impossible in theory for data assimilation to tell the difference between two solutions initialized of the form Equation˜29. That is, unlike the case of Figure˜5 if one considers two solutions $u_{1}$ and $u_{2}$ of Equation˜19 initialized as in Equation˜29 with separate $k_{0}$ values for $u_{1}$ and $u_{2}$, we have that $\mathinner{\!\left\lVert P_{M}(u_{1}-u_{2})\right\rVert}_{L^{2}}\neq 0$ for $t>0$, and to the best of our knowledge one cannot readily construct two such solutions of this form that always agree on every observed wavemode for all time. This is due to the fact that dispersive term essentially causes the coefficients of the wavemodes to rotate in the complex plane around the origin with rotation speed $\delta^{2}k^{3}$. As each wavemode is rotating at a different speed one cannot rescale and shift momentum across wavemodes in a straightforward way such that the momentum backscattering due to the nonlinearity on the observed modes still agree at each timestep to produce a counterexample similar to the one present in Figures˜5 and 4.
All of the cases where nudging fails are specific to the undamped and unforced, classical KdV Equation˜19, which is not a dissipative dynamical system. The addition of damping will result in the energy being drained from these high oscillations, while the forcing injects energy into specified low wavemodes, resulting in nontrivial long time dynamics that eliminate these sorts of extremely high frequency oscillations, see Section˜3.2.2 below.
@@ 3.2.2. Damped and Driven KdV
In [53] it was shown analytically, that given enough observed Fourier modes and large enough $\mu$, that the nudged solution converges to the reference solution in $L^{2}$ exponentially fast in time with exponent $\frac{\gamma}{4}$ for arbitrary choice of initial data $v^{in}\in H^{2}$ with spatial mean zero for Equation˜33.
@@
In Figure˜9 we can see nudging effectively recovers a reference solutions initialized as in Equation˜29 exponentially fast in time. Here we utilize $\mu=100$ and $\gamma=0.1$ and the exponential rate of convergence is given by small perturbations from $\gamma$. Here we initialize the reference solution as in Equation˜29 with $k_{0}=12$, that is, we selected the initial profile $u^{in}=\cos(12\cdot\pi\cdot x)$. For our driving force, we used $f=\cos(8\cdot\pi\cdot x)$. In the simulations shown in this section we use $M=10$, indicating that only the first 10 Fourier modes of the solution are observed. It is not surprising that convergence happens at the exponential rate proportional to the damping coefficient, $\gamma$, as this coincides with the results of [53]. We note that the exponential convergence is not entirely independent of $\mu$, the oscillatory component at initial times and on the observed modes is particularly sensitive to $\mu$ values, although this effect quickly fades as the error begins to decay at rate $\gamma$. Moreover, as one sees from the spectrum plots in Figure˜9 the primary barrier for convergence are the spikes in undetected frequencies, therefore it is not surprising that convergence to the reference solution happens at exactly the rate, $\gamma$, at which momentum is drained from these modes. Indeed, this behavior is typical of simulations of nudging methods, as the unobserved modes rely on the dissipation in the system in order to synchronize. The overall decay in error can typically be decomposed into two distinct phases, an initial rapid decay in error that is highly dependent on $\mu$, and a slower, but still exponentially fast in time, rate of convergence primarily driven by the dissipation in the system. For an in-depth study of the role of the particular value of $\mu$ see, e.g., [15].
@@
The AOT algorithm successfully recovers the reference solution, $u$, to Equation˜32 even when $u$ is initialized as in the counterexamples from the previous section Equations˜29 and 31 given enough observed modes. Note that in Figure˜9 we see some behavior in the observed modes due to the presence of the body force. The body force can be disabled by setting it to be identically zero, however in that case data assimilation would not be required at all. This is because the damping term will drive any initial condition to 0, exponentially fast at the rate $\gamma$ with or without any form of data assimilation being applied. Additionally, the AOT algorithm is only guaranteed to recover the true solution of Equation˜32 asymptotically in time, which in the case of zero body forcing is the zero solution. While the observed scales may be approximately recovered via the AOT algorithm, solutions initialized as in Equations˜29 and 31 will exhibit large oscillations in the unobserved Fourier modes that only approximately vanish after advancing suitably far in time.
@@ 4.1. Analytical Results
We now extend our results to the incompressible Euler equations on a 2D periodic domain. The equations are given as follows:
$\begin{split}\boldsymbol{u}_{t}+\boldsymbol{u}\cdot\nabla\boldsymbol{u}&=-\nabla p,\\ \divergence\boldsymbol{u}&=0.\end{split}$ (34)
@@
$\begin{split}\boldsymbol{v}_{t}+\boldsymbol{v}\cdot\nabla\boldsymbol{v}&=-\nabla q+\mu P_{M}(\boldsymbol{u}-\boldsymbol{v}),\\ \divergence\boldsymbol{v}&=0.\end{split}$ (35)
Here $\boldsymbol{v}(x,t)=(v_{1}(\boldsymbol{x},t),v_{2}(\boldsymbol{x},t))$ is the velocity field, $q$ is the pressure, $\mu$ is the nudging parameter, and $M$ is the number of observed modes in the reference solution $u$ of Equation˜34. As we are observing a 2D solution, $u$ of the system Equation˜34, $P_{M}$ denotes the observations of all Fourier modes $\boldsymbol{k}\in\mathbb{Z}^{2}\setminus\{(0,0)\}$ such that $\mathinner{\!\left\lvert\boldsymbol{k}\right\rvert}\leq M$.
@@
$(\boldsymbol{u},\boldsymbol{v})=\sum_{m=1}^{2}\int_{\Omega}u_{m}v_{m}d{\boldsymbol{x}}\quad\text{ and }\quad((\boldsymbol{u},\boldsymbol{v}))=\sum_{m,n=1}^{2}\int_{\Omega}\partialderivative{u_{m}}{x_{n}}\partialderivative{v_{m}}{x_{n}}d{\boldsymbol{x}},$ (37)
which are associated with the norms $\mathinner{\!\left\lVert\boldsymbol{u}\right\rVert}_{H}=\mathinner{\!\left\lVert\boldsymbol{u}\right\rVert}_{L^{2}}=(\boldsymbol{u},\boldsymbol{u})^{1/2}$ and $\mathinner{\!\left\lVert\boldsymbol{v}\right\rVert}_{V}=((\boldsymbol{u},\boldsymbol{u}))^{1/2}$. We note that the $V$ norm is the $H^{1}$ seminorm, which we use as the full norm for $V$ which is justified by the Poincaré inequality. We denote by $V^{s}$ the equivalent Sobolev space, given by $V^{s}=\overline{\mathcal{V}}^{V^{s}}$, where the $V^{s}$ norm is given by the $H^{s}$ seminorm.
We now apply the Leray-Helmholtz projector, $P_{\sigma}$, to Equations˜34 and 35 to formally obtain the following evolution equations for $\boldsymbol{u}$ and $\boldsymbol{v}$, respectively
@@
Here we formally denote the nonlinear operator, $B(\boldsymbol{a},\boldsymbol{b})\mathrel{\mathop{\ordinarycolon}}=P_{\sigma}(\boldsymbol{a}\cdot\gradient\boldsymbol{b})$ (cf. [76, 24]).
We recall the following theorem regarding global well-posedness of solutions to Equation˜34
Theorem 4.1 (Global well-posedness of Equation˜34 (see e.g. [64])).
Given $\boldsymbol{u}^{in}\in V^{s}(\mathbb{T}^{2})$ with $s\geq 3$, then Equation˜34 is globally well-posed with solution
@@
We now present the analogues of the results from Section˜3.2.1 for the 2D incompressible Euler equations. As the proofs follow similarly to their counterparts in Section˜3.2.1, we summarize them into the following key statements without detailed proofs.
@@
Let $k\in\mathbb{Z}\setminus\{0\}$ and let $\boldsymbol{u}^{in}\in V^{3}$ be given with zero spatial mean, satisfying for all $\boldsymbol{x}\in\mathbb{R}^{2}$ and each $j=1,2$,
$\boldsymbol{u}^{in}(\boldsymbol{x}+\frac{L}{k}\boldsymbol{e}_{j})=\boldsymbol{u}^{in}(\boldsymbol{x}).$
If $\boldsymbol{u}(\boldsymbol{x},t)$ is the solution of Equation˜34 with initial data $\boldsymbol{u}^{in}$, then for all $\boldsymbol{x}\in\mathbb{R}^{2}$, $t\in\mathbb{R}$, and each $j=1,2$, $\boldsymbol{u}(\boldsymbol{x},t)$ satisfies
$\boldsymbol{u}(\boldsymbol{x}+\frac{L}{k}\boldsymbol{e}_{j},t)=\boldsymbol{u}(\boldsymbol{x},t).$
In particular, the space of functions periodic with respect to the lattice $k\mathbb{Z}^{2}\setminus\{\boldsymbol{0}\}$ is invariant under the solution of Equation˜34.
@@
Let $k\in\mathbb{Z}\setminus\{0\}$ and let $\boldsymbol{u}^{in}_{j}\in V^{3}$, for $j=1,2$, satisfying the conditions of Proposition˜4.2. Moreover, assume $\mathinner{\!\left\lVert\boldsymbol{u}^{in}_{1}\right\rVert}_{H}\neq\mathinner{\!\left\lVert\boldsymbol{u}^{in}_{2}\right\rVert}_{H}$ and let $\boldsymbol{u}_{j}(x,t)$ be the corresponding solutions of Equation˜34 with initial data $\boldsymbol{u}_{j}^{in}$. Then for any $M<\mathinner{\!\left\lvert k\right\rvert}$ and all $t\geq 0$,
$P_{M}(\boldsymbol{u}_{1}(\cdot,t)-\boldsymbol{u}_{2}(\cdot,t))=0.$ (40)
and
$\limsup_{t\to\infty}\mathinner{\!\left\lVert\boldsymbol{u}_{1}(\cdot,t)-\boldsymbol{u}_{2}(\cdot,t)\right\rVert}_{H}>0.$ (41)
Now, it follows immediately from Proposition˜4.3 that solutions to Equation˜34 do not have the finitely many determining modes property:
Theorem 4.4.
The incompressible Euler equations Equation˜34 subject to periodic boundary conditions on $\mathbb{T}^{2}$ do not possess the finitely many determining modes property in $H$.
@@
For a fixed number of observed Fourier modes, $M$, there exists solutions $\boldsymbol{u}(\boldsymbol{x},t)$ of Equation˜34 that cannot be recovered by Equation˜35 with arbitrary initial condition $\boldsymbol{v}^{in}$.
In a future work, we will extend these results to the damped and driven Euler equations, which introduce additional complexities due to the external forcing and damping terms.